OpenStudy (erniellyd28):

a rental truck can be rented from company A for $110 a day plus $0.20 per mile. company B charges $70 a day plus $70 per mile to rent the same truck. how many miles must be driven in a day to make the rental cost for rental cost for company A a better deal than company b's

OpenStudy (aveline):

Can you write an equation?

OpenStudy (erniellyd28):

there were no directions to write one is that what you are asking me. (i'm very sorry for the question)

OpenStudy (aveline):

Ok. Are you sure that its $70 per mile?

OpenStudy (erniellyd28):

$0.70

OpenStudy (erniellyd28):

sorry lol

OpenStudy (aveline):

Ok. Let's use the equation: 110+0.20m=70+70m The 'm' represents the numbers of miles driven Let's simplify the equation: 110+0.20m=70+0.70m 40+0.20m=0.70m 40=0.50m 80=m So after 80 miles, the costs will be about the same. Therefore, after 81 miles...

OpenStudy (erniellyd28):

so 81 is the answer?

OpenStudy (aveline):

81 miles is the answer. It would cost $126.20 to rent from company A after 81 miles. It would cost $126.70 from company B.

OpenStudy (erniellyd28):

thank you very much for your help :-)